A thief is running away on a straight road in a jeep moving with a speed of 9 m /s . A policeman chases him on a motor cycle moving at a speed of 10 m/s . If the instantaneous separation of the jeep from the motor cycle is 100 m, how long will it take for the policeman to catch the thief? (A) 1 s (B) 19 s (C) 90 s (D) 100 s

Grade:

2 Answers

Saurabh Kumar
9 years ago
Time = Seperation between them /Relative velocity of approach.....= 100/(10-9) = 100 sec.
Vikas TU
9 years ago
Time = Seperation between them /Relative velocity of approach.....= 100/(10-9) = 100 sec.